OpenStudy (123ab456c):

What is the area in square inches of a rectangle with a length of 6.3 feet, and a width of 2.5 feet? A. 189 in2 B. 2268 in2 C. 141.75 in2 D. 15.75 in2

OpenStudy (123ab456c):

i think the answer is D

OpenStudy (retireed):

Convert feet to inches!!

OpenStudy (anitas.):

no A is incorrect!

OpenStudy (123ab456c):

really can you explain to me how to

OpenStudy (anitas.):

Area of a rectangle is L * W A = 6.3 * 2.5 A = 15.75 ft2 which, to this calculator converting sq ft to sq would be A = 2268 in2

OpenStudy (retireed):

6.3 ft times 2.5 feet = 15.75 Feet squared 15.75 Ft^2 * 144 inches^2 / 1 Ft^2 = _________ in^2 Ft^2 cancel

OpenStudy (123ab456c):

so it is B?

OpenStudy (anitas.):

yes because we multiply the lenght times width

OpenStudy (123ab456c):

thanks i fanned both of you!:):)